A 6.15 kg block is pulled by a rope up a ramp with an incline of 25 degrees from the horizontal. The acceleration of the block is 2.45 m/s^2. The coefficient of friction between the block and the ramp is 0.266. What is the tension in the rope?
Answer #1

Fnet = FT - m*g*sin(25) - μ*k*m*g*cos(25)

m*a = FT - m*g*sin(25) - μ*k*m*g*cos(25)

FT = m*(a + g*sin(25) + μ*k*g*cos(25))

= 6.15*(2.45 + 9.81*sin(25) + 0.266*9.81*cos(25))

= 55.11 N